# Berkovich Motives

### v1

@@ 5.2. The Tate twist
We work in the nonarchimedean locus $|2|\leq 1$. Then the strong triangle inequality holds on all uniform Banach rings, and in particular for any $r\in\mathbb{R}$, there is a subgroup $A_{<r}\subset A$ of elements of norm $<r$. Similarly, if $r\leq 1$, then $1+A_{<r}\subset A^{\times}$ is stable under multiplication. We call $1+A_{<1}\subset A^{\times}$ the group of $1$-units, and denote the corresponding sheaf by $1+\mathcal{O}_{<1}$.
@@
It remains to handle the situation after rationalization. In this case, there is no Galois cohomology. Thus, it reduces to a computation of the cohomology of the naively defined sheaf $\overline{\mathbb{G}}_{m}\otimes\mathbb{Q}$ (the sheafification of $A\mapsto A^{\times}/(1+A_{<1})\otimes\mathbb{Q}$) on the topological space $\mathcal{M}(C\langle T\rangle_{1})$. As this is a tree, there can only be cohomology in degree 00 and $1$. Moreover, by right-exactness, the vanishing of $H^{1}$ reduces to the vanishing of $H^{1}$ with coefficients in $\mathbb{G}_{m}$, which follows from $\mathrm{Pic}(C\langle T\rangle_{1})=0$. It remains to see that
@@
$\overline{\mathbb{G}}_{m}\otimes\mathbb{Q}\to\mathbb{R}_{>0}\xrightarrow{\mathrm{log}}\mathbb{R}.$
We claim that any global section on $\mathcal{M}(C\langle T\rangle_{1})$ maps to a constant. Indeed, the resulting map $\mathcal{M}(C\langle T\rangle_{1})\to\mathbb{R}$ factors locally over some quotient
@@
for some finite subset $A\subset\mathcal{O}_{C}$, and then by quasicompactness even globally. This gives a map from a tree to $\mathbb{R}$, and this map is piecewise-linear, and at each branch point except possibly the central Gaußpoint the sums of all slopes is equal to 00 – the slopes of the rays can locally be computed in terms of the valuations of elements in the residue fields of type (2) points, and the product formula for $\mathbb{P}^{1}$ then ensures the harmonicity. But on each ray $[0,1]$ of the tree, the map is constant near 00, and hence it must be constant on the whole tree.
Rescaling, we can assume that it is equal to 00. Looking at the Gauß point, or any type $2$ point, we get an element of $k(T)^{\times}\otimes\mathbb{Q}$, where $k$ is the residue field of $C$; but it must have trivial valuations, as the norm map is constant (and the valuations describe the derivatives along the rays). This works for all valuations except one valuation for the central Gauß point, but by the product formula, the valuation there is also trivial. Hence we get an element of $k^{\times}\otimes\mathbb{Q}$. At points of other types, the residue field is unchanged, and hence we necessarily get an element of $k^{\times}\otimes\mathbb{Q}$. By finitaryness, the resulting element of $k^{\times}\otimes\mathbb{Q}$ is locally constant, and as the space is connected, it is constant. ∎
@@
[…]
where the upper map is given by the characteristic polynomial and is an isomorphism of arc-sheaves. The left vertical and right vertical maps become equivalences on ball-invariant sheaves; thus, the lower map is an isomorphism on ball-invariant sheaves, which gives the desired result. ∎
@@
become isomorphisms on ball-invariant sheaves. But the projection to $\mathbb{G}_{m}$ yields a surjective map
@@
whose fibres are open balls (by a simple argument using Newton polygons); thus, on ball-invariant sheaves, the two are identified. In the colimit over $n$, the $\mathbb{T}(r_{1}^{n},r_{2}^{n})$ become all of $\mathbb{G}_{m}$.
In particular, the functor $\mathrm{Hom}(\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}],-)$ agrees with $\mathrm{Hom}(\mathbb{Z}_{\mathrm{mot}}[\mathbb{T}],-)$, i.e. evaluation at $\mathbb{T}$, for all annuli $\mathbb{T}$. Writing a closed annulus as an intersection of open annuli, and using finitaryness, it follows that it also agrees with evaluation at a(ny) closed annulus. But a closed annulus has finite cohomological dimension, so evaluation commutes with all colimits, so $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]$ and hence $\mathbb{Z}(1)$ is compact. ∎
@@ 6. Free motivic sheaves
In the last section, we proved the existence of free motivic sheaves $\mathbb{Z}_{\mathrm{mot}}[U]$ where $U$ is an open subspace of affine space $\mathbb{A}^{n}_{X}$. Our first goal is to make these more explicit. We will restrict attention to the nonarchimedean locus. In that case, we have seen that these free motivic sheaves can be understood in terms of those of $U/\mathbb{D}^{n}(0,<r)$ for $r>0$, which are even finitary over $X$.
@@
    After rationalization, this holds true also for the values on any affine subset of $\mathbb{A}^{1}_{A}$.
@@
Let $M$ be a finitary ball-invariant presheaf of abelian groups with transfer. Then the arc-sheafification of $M$, taken in $\mathcal{D}_{\mathrm{fin}}(X,\mathbb{Z})$, is also ball-invariant, after inverting all primes $p$ such that $|p|<1$ somewhere on $X$. After reduction modulo $n$ for some $n$ with $|n|\geq 1$ everywhere on $X$, it is invariant under change of algebraically closed field.
@@
The arc-sheafification is still finitary. The ball-invariance is a statement about values on algebraically closed nonarchimedean Banach fields $C$, so we can assume that $X=\mathcal{M}_{\mathrm{arc}}(C)$. We can consider separately the cases where $M$ is torsion, and where $M$ is rational. In the torsion case, we have to see that $M$ is invariant under change of algebraically closed field. Thus, let $C^{\prime}/C$ be an extension of algebraically closed Banach fields. By induction, we can assume that $C^{\prime}$ has topological transcendence degree $1$ over $C$. Then $\mathcal{M}_{\mathrm{arc}}(C^{\prime})$ can be written as a cofiltered limit of $U_{i}$ where $U_{i}/C$ are connected smooth Berkovich curves. By finitaryness, any $C^{\prime}$-point is induced from some $U_{i}$-point; as $U_{i}$ has $C$-points, this already implies injectivity of $M(C)\to M(C^{\prime})$. For surjectivity, we can replace $M$ by the image of $\mathbb{N}[U_{i}]\to M$. Applying Lemma 6.4 to the base change of $U_{i}$ to $C^{\prime}$ and the $C^{\prime}$-point of interest $x^{\prime}$, as well as any chosen $C$-point $x$ of $U_{i}$, we conclude.
@@
is a split short exact sequence in the category of ball-invariant presheaves of $\mathbb{Q}$-vector spaces with transfer. For this claim, which is purely diagrammatic, it is sufficient to understand the very small part of the category of ball-invariant presheaves of abelian groups with transfer that is spanned by the four objects
@@
Potentially after shrinking $U$ (still containing $x$ and $x^{\prime}$) we can find a smooth projective compactification $X$ of $U$ such that $Z=X\setminus U$ is a finite union of closed discs. Choose one point $z_{i}\in Z$ in each of those boundary discs, and consider the moduli space of degree 00 line bundles $\mathcal{L}$ on $X$ with trivializations at all $z_{i}$. This is a semiabelian variety, and in particular multiplication by $p$ is surjective. We can then divide $\mathcal{O}([x]-[x^{\prime}])$ by $p$ to get some line bundle $\mathcal{L}$, trivialized at the $z_{i}$, with $\mathcal{L}^{\otimes p}\cong\mathcal{O}([x]-[x^{\prime}])$, compatibly with the trivializations at the $z_{i}$. We can then find a rational section $s$ of $\mathcal{L}$ without zeroes or poles on the boundary discs, and equal to $1$ at the $z_{i}$. This gives an isomorphism $\mathcal{L}\cong\mathcal{O}(-D)$ for some divisor $D$ (necessarily of degree 00) supported on $U$, so we can write
@@
that is moreover the identity at the fibres at $z_{i}$. Let $s_{1}$ and $s_{2}$ be the corresponding tautological sections of this line bundle; both are equal to $1$ at $z_{i}$. Taking a variable $t\in\mathbb{B}$, we get the section $ts_{1}+(1-t)s_{2}$ which for $t=0$ is $s_{2}$ but for $t=1$ is $s_{1}$. The value at $z_{i}$ is constant $1$, and this implies that these rational sections have no zeroes or poles on the boundary discs, so their divisors are entirely supported on $U\subset X$. This gives the desired homotopy. ∎
@@
Let $C$ be a non-discrete algebraically closed nonarchimedean Banach field, and let $U,V\subset\mathbb{A}^{1}_{C}$ be two open subsets whose complements in $\mathbb{P}^{1}_{A}$ are finite disjoint unions of closed discs. Then the value of $\mathbb{Q}_{\mathrm{mot}}[U]$ on $V$ is concentrated in degree 00 and given by the cokernel of
@@
Assume that the complement of $U$ in $\mathbb{P}^{1}_{A}$ is the union of closed discs $Z_{1},\ldots,Z_{d}$. Then $\mathbb{Q}_{\mathrm{mot}}[U]\cong\mathbb{Q}\oplus\overline{\mathbb{G}}_{m}^{d}$. We may similarly describe $\mathbb{Q}_{\mathrm{mot}}[V]$. There are no maps $\overline{\mathbb{G}}_{m}\to\mathbb{Q}$ as those correspond to the reduced $\mathbb{Q}$-cohomology of $\mathbb{P}^{1}$ which vanishes. The endomorphisms of $\overline{\mathbb{G}}_{m}$ are given by $\mathbb{Q}$, concentrated in degree 00 (they can be computed using the value of $\overline{\mathbb{G}}_{m}$ at a type (3) point, which we may assume exists by reducing to the case that $C$ is separable). We see that the maps between $\mathbb{Q}_{\mathrm{mot}}[U]$ and $\mathbb{Q}_{\mathrm{mot}}[V]$ are concentrated in degree 00, and have a very explicit description. It is a pleasing exercise, similar to the proof of the previous lemma, to identify it with the displayed cokernel. The key observation is that given two maps $V\to\mathrm{Sym}^{n}(U)$, the agreement of the corresponding motivic invariants extracted above implies that the straight-line interpolation of the two induced maps $V\to\mathrm{Sym}^{n}(U)\subset\mathbb{A}^{n}_{C}$ (via the characteristic polynomial) still lies in $\mathrm{Sym}^{n}(U)$. This can be checked on each boundary disc individually and after replacing $V$ by a $C$-point, and then it is a simple observation about Newton polygons. ∎
@@
We have already seen that if $X$ is invariant under change of algebraically closed field, then $X$ is ball-invariant. For the converse, we have to see that for all open $U\subset\mathbb{A}^{m}_{X}$, the sheaf $\mathbb{Z}_{\mathrm{mot}}[U]/n$ is invariant under change of algebraically closed field. This follows from Theorem 6.1
@@
Let $p$ be a prime. If $|p|_{A}<1$, then $p$ is invertible in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$. If $|p|\geq 1$ everywhere on $\mathcal{M}(A)$, then the subcategory of $p$-power torsion objects on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$ is equivalent to the left-completed derived category of $p$-power torsion étale sheaves.
@@
For the first part, we can assume that $A$ is a Banach field. It is then perfectoid, and the arc-site is invariant under tilting; moreover, ball-invariance over $A$ implies invariance under the perfectoid ball by passing to filtered colimits, and hence ball-invariance on the tilt. (The converse is also true, see below.) Thus, we can assume that $A$ is of characteristic $p$. Here, the Artin-Schreier sequence $0\to\mathbb{F}_{p}\to\mathbb{B}\xrightarrow{x^{p}-x}\mathbb{B}\to 0$ shows that $\mathbb{F}_{p}=0$ in ball-invariant sheaves.
Now assume that $|p|\geq 1$ everywhere on $\mathcal{M}(A)$. By descent, we can reduce to the case that $X$ is strictly totally disconnected, in which case the subcategory of $p$-power torsion objects of $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$ is equivalent to $p$-power torsion sheaves on the profinite set $\mathcal{M}(A)$, by the previous corollary. Those are also the étale sheaves, as desired. ∎

### v2

@@ 5.2. The Tate twist
We work in the nonarchimedean locus $|2|\leq 1$. Then the strong triangle inequality holds on all uniform Banach rings, and in particular for any $r\in\mathbb{R}$, there is a subgroup $A_{<r}\subset A$ of elements of norm <r<r. Similarly, if $r\leq 1$, then $1+A_{<r}\subset A^{\times}$ is stable under multiplication. We call $1+A_{<1}\subset A^{\times}$ the group of $1$-units, and denote the corresponding sheaf by $1+\mathcal{O}_{<1}$.
@@
It remains to handle the situation after rationalization. In this case, there is no Galois cohomology. Thus, it reduces to a computation of the cohomology of the naively defined sheaf $\overline{\mathbb{G}}_{m}\otimes\mathbb{Q}$ (the sheafification of $A\mapsto A^{\times}/(1+A_{<1})\otimes\mathbb{Q}$) on the topological space $\mathcal{M}(C\langle T\rangle_{1})$. As this is a tree, there can only be cohomology in degree 0 and $1$. Moreover, by right-exactness, the vanishing of $H^{1}$ reduces to the vanishing of $H^{1}$ with coefficients in $\mathbb{G}_{m}$. Any class in this $H^{1}$ can be realized by a Čech $1$-cocycle, which can be used to build an invertible module over $C\langle T\rangle_{1}$ using analytic descent of line bundles on rigid-analytic spaces. Now $\mathrm{Pic}(C\langle T\rangle_{1})=0$ as $C\langle T\rangle_{1}$ is a principal ideal domain. Picking a global invertible section of this line bundle then trivializes the Čech cocycle, giving the vanishing of $H^{1}$. It remains to see that
@@
$\overline{\mathbb{G}}_{m}\otimes\mathbb{Q}\to\mathbb{R}_{>0}\xrightarrow{\mathrm{log}}\mathbb{R}$
of arc-sheaves (in particular, of sheaves on $\mathcal{M}(C\langle T\rangle_{1})$), where $\mathbb{R}$ denotes the arc-sheaf taking any $A$ to the set of continuous maps from $\mathcal{M}(A)$ to $\mathbb{R}$. We claim that any global section on $\mathcal{M}(C\langle T\rangle_{1})$ maps to a constant. Indeed, representing a section of $\overline{\mathbb{G}}_{m}$ locally by a section of $\mathbb{G}_{m}$, and approximating these invertible functions by a function of the form $\prod_{i}(T-a_{i})^{k_{i}}$, the resulting map $\mathcal{M}(C\langle T\rangle_{1})\to\mathbb{R}$ factors locally over some quotient
@@
for some finite subset $A\subset\mathcal{O}_{C}$, and then by quasicompactness even globally. This gives a map from a tree to $\mathbb{R}$, and this map is piecewise a linear combination of $\mathrm{log}(x_{a})=\mathrm{log}|T-a|$, and at each branch point except possibly the central Gauß point the sums of all slopes is equal to 0 – the slopes of the rays can locally be computed in terms of the valuations of elements in the residue fields of type (2) points, and the product formula for $\mathbb{P}^{1}$ then ensures the harmonicity. But on each ray $[0,1]$ of the tree, the map is constant near 0, and hence it must be constant on the whole tree.
Rescaling, we can assume that it is equal to 0. Looking at the Gauß point, or any type (2) point, we get an element of $k(T)^{\times}\otimes\mathbb{Q}$, where $k$ is the residue field of $C$; but it must have trivial valuations, as the norm map is constant (and the valuations describe the derivatives along the rays). This works for all valuations except one valuation for the central Gauß point, but by the product formula, the valuation there is also trivial. Hence we get an element of $k^{\times}\otimes\mathbb{Q}$. At points of other types, the residue field is unchanged, and hence we necessarily get an element of $k^{\times}\otimes\mathbb{Q}$. By finitaryness, the resulting element of $k^{\times}\otimes\mathbb{Q}$ is locally constant, and as the space is connected, it is constant. ∎
@@
${\mathrm{Sym}^{n}(\mathbb{G}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{A}^{n-1}\times\mathbb{G}_{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sym}^{n}(\overline{\mathbb{G}}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{\mathbb{G}}_{m}}$
where the upper map is given by the characteristic polynomial and is an isomorphism of arc-sheaves. The left vertical and right vertical maps are (colimits of) $\mathbb{B}$-homotopy equivalences (for the left vertical map, write $\overline{\mathbb{G}}_{m}$ as the geometric realization $\mathbb{G}_{m}\times(1+\mathcal{O}_{<1})^{\bullet}$, all of whose terms are $\mathbb{B}$-homotopy equivalent to $\mathbb{G}_{m}$; then apply $\mathrm{Sym}^{n}$ everywhere); thus, the lower map is an isomorphism when mapping to ball-invariant sheaves, which gives the desired result. ∎
@@
become isomorphisms when mapping into ball-invariant sheaves. But the projection to $\mathbb{G}_{m}$ yields a surjective map
@@
whose fibres are open balls: Indeed, a polynomial
$X^{n}+a_{n-1}X^{n-1}+\ldots+a_{1}X+a_{0}$
with given $a_{0}\in\mathbb{T}(r_{1}^{n},r_{2}^{n})$ has all zeros in $\mathbb{T}$ if and only if all slopes of the Newton polygon defined by $a_{n-1},\ldots,a_{1},a_{0}$ lie in the interval $(r_{1},r_{2})$. This means
$|a_{i}|<\mathrm{max}(r_{2}^{n-i},|a_{0}|r_{1}^{-i})$
for all $i=1,\ldots,n-1$, which is indeed an open ball. Thus, when mapping to ball-invariant sheaves, the map $\mathrm{Sym}^{n}(\mathbb{T})\to\mathbb{T}(r_{1}^{n},r_{2}^{n})$ becomes an isomorphism. In the colimit over $n$, the $\mathbb{T}(r_{1}^{n},r_{2}^{n})$ become all of $\mathbb{G}_{m}$.
In particular, the functor $\mathrm{Hom}(\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}],-)$ agrees with $\mathrm{Hom}(\mathbb{Z}_{\mathrm{mot}}[\mathbb{T}],-)$, i.e. evaluation at $\mathbb{T}$, for all annuli $\mathbb{T}$. Writing a closed annulus as an intersection of open annuli, and using finitaryness, it follows that it also agrees with evaluation at a(ny) closed annulus. But a closed annulus has finite cohomological dimension and is quasicompact, so evaluation commutes with all colimits, so $\mathbb{Z}_{\mathrm{mot}}[\mathbb{G}_{m}]$ and hence $\mathbb{Z}(1)$ is compact. ∎
@@ 6. Free motivic sheaves
In the last section, we proved the existence of free motivic sheaves $\mathbb{Z}_{\mathrm{mot}}[U]$ where $U$ is an open subspace of affine space $\mathbb{A}^{n}_{X}$. Our first goal is to make these more explicit. We will restrict attention to the nonarchimedean locus. In that case, we have seen that these free motivic sheaves can be understood in terms of those of $U/\mathbb{D}^{n}(0,<r)$ for some $r>0$, which are even finitary over $X$.
@@
    After rationalization, this holds true also for the values on any affine subset of $\mathbb{A}^{1}_{A}$, i.e. any arc-subsheaf of the form $\mathcal{M}_{\mathrm{arc}}(B)$ for some Banach $A$-algebra $B$.
@@
Let $M$ be a finitary ball-invariant presheaf of abelian groups with transfer over a strictly totally disconnected $X$. Then the derived arc-sheafification of $M$ is also ball-invariant, after inverting all primes $p$ such that $|p|<1$ somewhere on $X$. After reduction modulo $n$ for some $n\geq 1$, it is invariant under change of algebraically closed field.
@@
The derived arc-sheafification can be computed by first restricting to strictly totally disconnected Banach rings $A$ over $X$, and then using Theorem 4.1 and pro-étale sheafification. This is still finitary by Lemma 4.7. The ball-invariance is a statement about values on algebraically closed nonarchimedean Banach fields $C$, so we can assume that $X=\mathcal{M}_{\mathrm{arc}}(C)$. We can consider separately the cases where $M$ is torsion, and where $M$ is rational. In the torsion case, we have to see that $M$ is invariant under change of algebraically closed field. Thus, let $C^{\prime}/C$ be an extension of algebraically closed Banach fields. By induction, we can assume that $C^{\prime}$ has topological transcendence degree $1$ over $C$. Then $\mathcal{M}_{\mathrm{arc}}(C^{\prime})$ can be written as a cofiltered limit of $U_{i}$ where $U_{i}/C$ are connected smooth Berkovich curves. By finitaryness, any $C^{\prime}$-point is induced from some $U_{i}$-point; as $U_{i}$ has a $C$-point $x\in U_{i}$, this already implies injectivity of $M(C)\to M(C^{\prime})$. For surjectivity, consider the $C$-point $x\in U_{i}$ just chosen as well as the tautological $C^{\prime}$-point $x^{\prime}$ of $U_{i}$. It suffices to see that the map $M(U_{i})\to M(C^{\prime})$ induced by $x^{\prime}$ agrees with the composite map $M(U_{i})\to M(C)\hookrightarrow M(C^{\prime})$ induced by $x$. This follows from Lemma 6.4.
@@
is a split short exact sequence in the category of ball-invariant presheaves of $\mathbb{Q}$-vector spaces with transfer. For this claim, which amounts only to the existence of certain maps satisfying certain identities, it is sufficient to understand the very small part of the category of ball-invariant presheaves of abelian groups with transfer that is spanned by the four objects
@@
Potentially after shrinking $U$ (still containing $x$ and $x^{\prime}$) we can find a smooth projective compactification $X$ of $U$ such that $Z=X\setminus U$ is a finite union of closed discs, see [vdP80, Corollary 1.3]. Choose one point $z_{i}\in Z$ in each of those boundary discs, and consider the moduli space of degree 0 line bundles $\mathcal{L}$ on $X$ with trivializations at all $z_{i}$. This is a semiabelian variety, and in particular multiplication by $p$ is surjective. We can then divide $\mathcal{O}([x]-[x^{\prime}])$ by $p$ to get some line bundle $\mathcal{L}$, trivialized at the $z_{i}$, with $\mathcal{L}^{\otimes p}\cong\mathcal{O}([x]-[x^{\prime}])$, compatibly with the trivializations at the $z_{i}$. We can then find a rational section $s$ of $\mathcal{L}$ without zeroes or poles on the boundary discs, and equal to $1$ at the $z_{i}$. This gives an isomorphism $\mathcal{L}\cong\mathcal{O}(-D)$ for some divisor $D$ (necessarily of degree 0) supported on $U$, so we can write
@@
that is moreover the identity at the fibres at $z_{i}$. Let $s_{1}$ and $s_{2}$ be the corresponding tautological sections of this line bundle; both are equal to $1$ at $z_{i}$. Taking a variable $t\in\mathbb{B}$, we get the section $ts_{1}+(1-t)s_{2}$ which for $t=0$ is $s_{2}$ but for $t=1$ is $s_{1}$. The value at $z_{i}$ is constant $1$, and this implies that these sections have no zeroes on the boundary discs (as $\frac{s_{1}}{s_{2}}$ is an invertible function on the disc and equal to $1$ at $z_{i}$, thus $|\frac{s_{1}}{s_{2}}-1|<1$ and therefore $1+t(\frac{s_{1}}{s_{2}}-1)=\frac{ts_{1}+(1-t)s_{2}}{s_{2}}$ is a $1$-unit), so their divisors are entirely supported on $U\subset X$. This gives the desired homotopy. ∎
@@
Let $C$ be a non-discrete algebraically closed nonarchimedean Banach field, and let $U,V\subset\mathbb{A}^{1}_{C}$ be two open subsets whose complements in $\mathbb{P}^{1}_{A}$ are finite disjoint unions of closed discs. Then the value of $\mathbb{Q}_{\mathrm{mot}}[U]$ on $V$ is concentrated in degree 0 and given by the cokernel of
@@
Assume that the complement of $U$ in $\mathbb{P}^{1}_{A}$ is the union of closed discs $Z_{0},\ldots,Z_{d}$, where $Z_{0}$ is a disc around $\infty$. Then $\mathbb{Q}_{\mathrm{mot}}[U]\cong\mathbb{Q}\oplus\overline{\mathbb{G}}_{m}^{d}$. We may similarly describe $\mathbb{Q}_{\mathrm{mot}}[V]$, invoking $e$ finite boundary discs. There are no maps $\overline{\mathbb{G}}_{m}\to\mathbb{Q}$ as those correspond to the reduced $\mathbb{Q}$-cohomology of $\mathbb{P}^{1}$ which vanishes. The endomorphisms of $\overline{\mathbb{G}}_{m}\otimes\mathbb{Q}$ are given by $\mathbb{Q}$, concentrated in degree 0 (they can be computed using the value of $\overline{\mathbb{G}}_{m}$ at a type (3) point, which we may assume exists by reducing to the case that $C$ is separable). We see that the maps between $\mathbb{Q}_{\mathrm{mot}}[U]$ and $\mathbb{Q}_{\mathrm{mot}}[V]$ are concentrated in degree 0, and have a very explicit description as
$\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{de}.$
Concretely, the first two components are equivalent to a map $\mathbb{Q}\to\mathbb{Q}_{\mathrm{mot}}[U]$, corresponding to restriction along a chosen point $v\in V$; and the component $\mathbb{Q}^{de}$ measures degrees of maps on boundary tori.
Now we first prove surjectivity of the map
$(\mathbb{N}[U])(V)\otimes\mathbb{Q}\to\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{de}.$
It suffices to do this in case $e=1$: In general write $V=V_{1}\cap\ldots\cap V_{e}$ and use the commutative diagram
${\bigoplus_{i}(\mathbb{N}[U])(V_{i})\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i}(\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{e})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\mathbb{N}[U])(V)\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{de}}$
where the right-vertical map is surjective. Thus, shifting and rescaling $V$ if necessary, we can assume $V=\mathbb{T}(s_{1},s_{2})$ for some $0<s_{1}<1<s_{2}$. Similarly, we can assume that
$U=\mathbb{D}(0,<r_{2})\setminus\bigcup_{i=1}^{d}\mathbb{D}(a_{i},\leq r_{1})$
for some $r_{2}>1>r_{1}$ and $a_{i}\in\mathcal{O}_{C}$, pairwise distinct in the residue field $k$. In general, there is an open immersion $\mathrm{Sym}^{n}(U)\subset\mathbb{A}^{n}_{C}$ given by the characteristic polynomial. We now claim that for any sufficiently large integers $k_{1},\ldots,k_{d}\geq 1$ the map
$V\to\mathbb{A}^{n}_{C}:t\mapsto\prod_{i=1}^{d}(T-a_{i})^{k_{i}}-t$
factors over $\mathrm{Sym}^{n}(U)$, where $n=k_{1}+\ldots+k_{d}$. Concretely, this means that any $T\in C$ satisfying
$\prod_{i=1}^{d}(T-a_{i})^{k_{i}}=t$
must lie in $U$. Now if we had $|T|\geq r_{1}>1$, then as all $|a_{i}|\leq 1$, the left-hand side has absolute value $r_{2}^{n}$, which is $s_{2}>|t|$ once $n$ is sufficiently large. On the other hand, if for some $i$ we have $|T-a_{i}|\leq r_{1}$, then for all other $j$ we have $|T-a_{i}|=|a_{i}|\leq 1$, and the absolute value of the left-hand side is at most $r_{1}^{n}$, which is $s_{1}<|t|$ once $n$ is sufficiently large. Thus indeed, the map factors over $\mathrm{Sym}^{n}(U)$. This means that the composite map
$(\mathbb{N}[U])(V)\otimes\mathbb{Q}\to\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{d}\to\mathbb{Q}^{d}$
has $(k_{1},\ldots,k_{d})$ in its image, for all sufficiently large $k_{1},\ldots,k_{d}$. This shows that this composite map is surjective. Now it suffices to see that the map
$(\mathbb{N}[U])(\ast)\otimes\mathbb{Q}\to\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}$
is surjective. But this takes an element $T$ of $U(C)$ to $(1,(T-a_{i})_{i})$. Pick two elements $T_{1},T_{2}$ with $|T_{1}-a_{i}|,|T_{2}-a_{i}|<1$ for some $i$. Then for $j\neq i$, $\frac{T_{1}-a_{j}}{T_{2}-a_{j}}=1+\frac{T_{2}-T_{1}}{T_{2}-a_{j}}$ is a $1$-unit, so the difference of the images of $T_{1}$ and $T_{2}$ is nonzero only in the $i$-th component, where it is the image of $\frac{T_{1}-a_{i}}{T_{2}-a_{i}}$. This shows that
$0\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\subset\mathbb{Q}\to\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}$
is in the image, but then the image is everything.
Finally, we need to prove injectivity. Thus consider two maps $V\to\mathrm{Sym}^{n}(U)$ yielding the same element of
$\mathbb{Q}\oplus(C^{\times}/(1+C_{<1}))^{d}\otimes\mathbb{Q}\oplus\mathbb{Q}^{de}.$
We regard them as maps $f_{0},f_{1}:V\to\mathrm{Sym}^{n}(U)\subset\mathbb{A}^{n}_{C}$, and interpolate to a family of maps
$(f_{t})_{t\in\mathbb{B}}:V\times\mathbb{B}\to\mathbb{A}^{n}_{C}$
via $f_{t}=(1-t)f_{0}+tf_{1}$. We claim that this factors over $\mathrm{Sym}^{n}(U)\subset\mathbb{A}^{n}_{C}$. This can be checked pointwise, so we reduce to the case that $V$ is a point. Writing $U$ as an intersection (and translating it), we can also assume that $U=\mathbb{T}(r_{1},r_{2})$ is an annulus. Then it follows from the observation, in the proof of Proposition 5.19, that
$\mathrm{Sym}^{n}(U)\to\mathbb{T}(r_{1}^{n},r_{2}^{n})$
is a fibration in open balls. Here, the assumption that $f_{0}$ and $f_{1}$ induce the same map on $\mathbb{Q}_{\mathrm{mot}}[-]$ ensures that they yield the same element of $\mathbb{T}(r_{1}^{n},r_{2}^{n})$ up to $1$-units. ∎
@@
We have already seen that if $M$ is invariant under change of algebraically closed field, then $M$ is ball-invariant. For the converse, we can assume that $X$ is strictly totally disconnected. Using Proposition 5.13, it suffices to see that for all open $U\subset\mathbb{A}^{m}_{X}$, the sheaf $\mathbb{Z}_{\mathrm{mot}}[U]/n$ is invariant under change of algebraically closed field. This follows from Theorem 6.1. ∎
@@
Let $p$ be a prime and let $A$ be a Banach ring. If $|p|_{A}<1$, then $p$ is invertible in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$. If $|p|\geq 1$ everywhere on $\mathcal{M}(A)$, then the subcategory of $p$-power torsion objects on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$ is equivalent to the left-completed derived category of $p$-power torsion étale sheaves.
@@
For the first part, we can assume that $A$ is an algebraically closed Banach field. It is then perfectoid, and the arc-site is invariant under tilting, cf. [Sch17, Theorem 3.12, 3.13]; moreover, ball-invariance over $A$ implies invariance under the perfectoid ball by passing to filtered colimits, and hence ball-invariance on the tilt. (The converse is also true, see below.) Thus, we can assume that $A$ is of characteristic $p$. Here, the Artin-Schreier sequence $0\to\mathbb{F}_{p}\to\mathbb{B}\xrightarrow{x^{p}-x}\mathbb{B}\to 0$ shows that $\mathbb{F}_{p}=0$ in ball-invariant sheaves.
Now assume that $|p|\geq 1$ everywhere on $\mathcal{M}(A)$. By descent, we can reduce to the case that $A$ is strictly totally disconnected, in which case the subcategory of $p$-power torsion objects of $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathcal{M}(A))$ is equivalent to $p$-power torsion sheaves on the profinite set $\mathcal{M}(A)$, by the previous corollary. Those are also the étale sheaves, as desired. ∎